OpenStudy (lena772):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7 B right?

hartnn (hartnn):

your attempt ?

OpenStudy (lena772):

Slope same for parallel lines

hartnn (hartnn):

your answer is correct, so i will assume you did it correctly :)

OpenStudy (lena772):

6=-3*-1+b 6=3+b 6-3=b 3=b y=-3x+3

hartnn (hartnn):

and there you proved it :)